Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
takaaki.yokozawa - 14:50 Friday 19 November 2021 (18952) Print this report
Trial to detect GR beam at TMSY PD
[mTakahashi, Yokozawa with remote Ushiba]

Finally we detected the green beam at TMSY PD and gigE camera.

1. We moved the PDA and its pedestal at the EYA chamber
(Fig.1. Current position may affect to PCal team, we will discuss later)

2. We removed the aluminum foil and the metal board at the +Y side of EYA chamber

3. mTakahashi-san generated temporal duct to prevent from dust

4. Yokozawa turned on the PD for green at TMSY and turned off the LEDs at the TMSY area

5. Ushiba-san tried to survey by oscillating the SR3 mirror and finally we found the green beam at the PD.

6. When we changed the exposure of the gigE camera for TMSY_G, we can see the image
Images attached to this report
Comments to this report:
takafumi.ushiba - 16:21 Friday 19 November 2021 (18957) Print this report

Additional information of GRY alignment.

Figure1 shows the SR3 OpLev signals (pitch and yaw) and TMSY PDA1 signal.
At around -17s, I closed GRY shutter to confirm the PD signals are really coming from Green beam.
As it can be seen, the PD signal changes when shutter was closed, which means the signal is realy coming from green beam.

Then, I changed the SR3 OpLev setpoint to (P,Y)=(-2.2,1.2).

Search Help